OpenStudy (bekkah323):

write the slope-intercept form of the equation of the line satisfying the given conditions. Through (-2,4); slope -3/4

OpenStudy (akashdeepdeb):

Point-Slope form: (y-y1) = m(x-x1) (x1,y1) = (-2,4) m = -3/4 Get the equation of the line! Then write that in the form of y =mx+b Okay?